OpenStudy (anonymous):

consider the line whose questionis 3x+y-2=0. then the slope of any line that is parallel to this line is ___?. the slope of any line that is perpendicular to this line is __.

OpenStudy (a1234):

Parallel lines have the same slope, but different y-intercepts. 3x+y-2=0 3x + y = 2 y = -3x + 2 The slope is -3.

OpenStudy (anonymous):

the slope of the parallel or perpendicular? its asking for both.

OpenStudy (anonymous):

to get the slope of a perp line simply take the slope of that line, -3 change the numerator and denominator -1/3 and then change the sign 1/3 so the answer is 1/3

OpenStudy (anonymous):

or mathematically speaking (slope of a line)(slope of the perp line for that line) = -1

OpenStudy (anonymous):

ok so it says " the slopeof any line that is perpendicular to this line__?

OpenStudy (a1234):

Perpendicular lines have their slope producing a product of -1. The one I showed above is for parallel. 1/3*-3/1 = -3/3 = -1 Yes, 1/3.

OpenStudy (anonymous):

so -3 for the parallel and -1 for perp?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

why you just want to copy the answer instead of trying to read and understand what we are typing

OpenStudy (anonymous):

we are spending our valuable time to help you understand something and you don't even care? this is not how this site works

OpenStudy (anonymous):

I am trying to understand..

OpenStudy (anonymous):

i do cARE THAT IS WHY I AM TRYING TO UNDERSTAND '

OpenStudy (whpalmer4):

Okay, here's my best effort at helping you understand. Original line is \[3x+y-2=0\]If we solve that for \(y\), we put it in slope-intercept form, where it is easy to see the slope \((m)\): \[3x+y-2=0\]\[3x+y=2\]\[y=-3x+2\]Compare that with the slope-intercept form:\[y=mx+b\]\[m = -3\]so our slope is \(-3\) and any parallel line to this line will have a slope of \(-3\). Perpendicular lines (with one notable exception) have the property that the product of their slopes \(= -1\). Another name for this relationship is that they are negative reciprocals. If the slope of the first line is \(m_1\), then the slope of the second line is \[m_2 = -\frac{1}{m_1}\]Any line perpendicular to our line with slope \(m = -3\) will have slope\[m = -\frac{1}{-3} = \frac{1}{3}\] I said there was a notable exception to this relationship. A vertical line \((x=c)\) has an undefined slope, and a horizontal line \((y=c)\) has a slope of 0. (\(c\) is a constant). To construct the perpendicular line to a vertical or horizontal line, you have to write a new equation by hand in the appropriate form. You can't calculate the new slope by taking a negative reciprocal.